OpenStudy (anonymous):

in a simple electric circuit, the current in resistor is measured as (2.50+_0.05)mA. the resistor is marked as having a value of 4.7ohm+_2 % if these values were used to calculate the power dissipated in the in the resistor . what would be the percentage uncertainty in the value.

OpenStudy (anonymous):

Hint: P=VI=IR*I=I^2R

OpenStudy (anonymous):

now, Find percentage uncertanity in current

OpenStudy (anonymous):

can u?

OpenStudy (anonymous):

trying ...

OpenStudy (anonymous):

Ok. let me help: (0.05/2.50 * 100)%=

OpenStudy (anonymous):

thats 2%..... right?

OpenStudy (anonymous):

got it ..

OpenStudy (anonymous):

Now, Total percentage error in the power calculated = 2* 2% + 2%=6%

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

yes ..i did it ... :)

OpenStudy (anonymous):

Great

OpenStudy (anonymous):

Did u understand? what I did

OpenStudy (anonymous):

yes multiplied the uncertainty which we got by solving with the uncertainty of value obtained ...right?

OpenStudy (anonymous):

Its actually because 

OpenStudy (anonymous):

So, percentage uncertainity in Power= 2 times the percentage uncertainity in Current + percentage uncertainity in resistance

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Do u have any other problem?

OpenStudy (anonymous):

yes wait

OpenStudy (anonymous):

First close this question